An object is dropped from rest from three different heights a) 1m b)1.2m c)1.5m

Calculate the speed of the object using all three heights just before it hits the ground. Then derive the speed of the object as a function of height dropped.

Please give proper reasoning and details. I will give thumbs up.
Answer #1

Know that a = vou doc a=g vdu ch Yww.com w 529H = (funcation of heigh) (نا when - im NG - 12x9 12x 9.881 - 519.6 VA = 4.42 miwhen H = 1.5 (iii) NA 2x9-8X1.5 f = 5.422 mios